OpenStudy (anonymous):

Use the laws of exponents to simplify (10^7/4)^2/5

OpenStudy (anonymous):

\[(10^{7/4)^{2/5}}\]

OpenStudy (lalaly):

10^ ( (7/4)*(2/5)) = 10^ 14/20 = 10^7/5

OpenStudy (lalaly):

sorry 10^7/10

OpenStudy (anonymous):

\[10^{7/10} correct\]

OpenStudy (lalaly):

:D

OpenStudy (anonymous):

you are good at this

OpenStudy (lalaly):

lol am here to help